# Undergraduate Course Allocation

### v1

@@ C.4 Calibration
We use the Random Serial Dictatorship with adjusted course reserves to calculate simulated moments and find $\pi^{*}$ that minimize a criterion function:
$||m-m^{s}(\pi)||_{W}^{2}=(m-m^{s}(\pi))^{\prime}W(m-m^{s}(\pi)),$ (C5)
where $m$ is a set of moments constructed from the empirical data, $m^{s}(\pi)$ is the average of moments constructed from $100$ simulation rounds for a given vector of parameters $\pi$, and $W$ is a weight matrix. We draw random components of student utility and student choice sets for each simulation round. We use the following procedure in our simulations (see Jalali, Rahmandad and Ghoddusi, 2015):
@@
  • Second Stage. The second stage of the calibration process is iterative. We start with the optimal weight matrix $W^{*}_{\ell}=W^{*}$ for $\ell=1$. We then use the optimal weight matrix to obtain more precise estimates for parameters $\pi^{*}_{\ell}$. Using the more precise estimate $\pi^{*}_{\ell}$, we compute the optimal weight matrix $W^{*}_{\ell+1}$ and repeat the process $\ell=2,3,...$ until the maximum change of calibrated parameters $|\pi^{*}_{l+1}-\pi^{*}_{l}|$ becomes smaller than $0.01$ to ensure convergence.
@@
We observe that the horizontal components are positive for students in their first and second years and mostly negative in their third and fourth years. This observation suggests that students prefer enrolling in more courses during their initial years of study. The tendency to enroll in fewer than five courses in the later years may be due to several factors. Such students might choose not full-term courses that are not in our data or have already accumulated enough credits for their degree, allowing them to ease their course load in the final years.
The middle portion of the table presents interaction terms of matrix $\Gamma$. Each non-zero interaction coefficient $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$ corresponds to the increase in student’s utility from taking a course in college $a^{\prime}$ (column) compared to a course in student’s college $a$ (row) with the same vertical component. Note that we normalize $\gamma_{aa}=0,a\in{\cal A}$. The general trend of negative utility changes emphasizes a preference for internal college courses, but specific positive changes indicate potential areas for beneficial inter-college course offerings. For example, students from colleges $C$ and $D$ value highly courses in college $A$.
The bottom portion of the table presents quantiles of $756$ coefficients of vertical component $z$, with each coefficient $z_{c}$ being responsible for the average popularity of the course among students from the same college. The values range from a minimum of $-2.32$ to a maximum of $2.28$. This wide range indicates significant variation in the popularity of courses within colleges. Most of the quantile values are negative, up to the $90th$ percentile, indicating that a majority of the courses have negative values. This result is most likely manifested by the student utility model that relies on college-year parameters to identify student preferences. Students randomly draw $80$ courses in their choice set but no more than $5$ courses. Our model, which relies on college-level and year-level parameters, matches this empirical pattern by putting the values of the vertical components of most courses in the negative region.
@@ Appendix D Additional Simulations
This appendix provides additional simulation results. First, for the simulations conducted in the main text, we present statistics on the equilibrium prices in the Pseudo-Market with Priorities (PMP) mechanism (Section D.1) and compare performance with some alternative mechanisms (Section D.2). Second, we supplement the main text with simulation results with a department-first priority structure (Section D.3). Last, we report robustness checks by considering several sizes for student choice sets and the standard deviation of the noise utility parameter (Section D.4).
D.1 The Prices in Pseudo-Market with Priority Equilibria
First, we present statistics on the equilibrium prices in the PMP mechanism for the simulations reported in Section 3.3. The top part of Table D.1 provides information about where the cutoff levels of priority are across courses and how high the prices are for students at the cutoff level of priority. The bottom part provides information about the percentage of students allocated different numbers of courses with positive prices within each year of study.
$70.8\%$ of courses have their cutoff level of priority at $1$. The large share of courses at the lowest priority level reflects the non-binding enrollment of most courses in the university data, where just $11.2\%$ of courses are at or above their maximum enrollment capacity (see Table 2 on p. 2). Above level $1$, there is a noticeable difference in the percentage of courses with odd and even cutoff levels of priority. Substantially more courses have odd cutoff levels ($8.0\%$, $8.5\%$, $5.9\%$ for levels $3$, $5$, and $7$, respectively) than even ones ($2.1\%$, $2.6\%$, $1.7\%$, $0.5\%$ for levels $2$, $4$, $6$, and $8$). The hybrid priority structure we utilize places students at odd priority levels in a course if they are ineligible for any of its reserved seats. The students eligible for a course’s reserved seats typically only comprise a small subset of the interested students. Furthermore, some courses do not reserve any seats. As a result, the cutoff level of priority often ends up at an odd level rather than an even one.
@@ D.1 The Prices in Pseudo-Market with Priority Equilibria
Cutoff level of priority $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
@@
Notes: This table presents statistics on the equilibrium prices in the Pseudo-Market with Priorities mechanism for the simulations reported in Section 3.3. Results are averages across $100$ runs with different random component draws.
Among courses with a cutoff level of priority of $1$, the average price at the cutoff is just $0.06$. As many courses are free for all students, the average cutoff price for these courses is considerably lower than for courses with a higher cutoff level. For higher cutoff levels, the average cutoff prices lie between $0.75$ and $0.94$, with higher prices for odd cutoff levels. Student budgets are assigned in $[1,1.25]$, meaning that, at the average cutoff price, each student can afford one but not two courses for which they are at the cutoff level of priority. Still, students manage to take $k=5$ courses by taking courses for which they are above the cutoff level of priority, where they face a priority-specific price of 00.
The bottom part of Table D.1 further explores the number of courses students pay a positive price for. Across all years of study, at least $36\%$ of students obtain all their course seats for free, and fewer than $13\%$ of students pay a positive price for more than one course. Many courses have a zero price for all students, leading some students to obtain their most preferred course schedule for free. In addition, the average price of a course with a cutoff level of priority above $1$ is more than half of the maximum student budget. So, students often cannot afford a second course with a positive price after obtaining one.
These findings differ across years of study. Less than $7\%$ of cutoff levels of priority are $7$ or $8$, implying that Year $4$ students face zero price for the vast majority of courses. Almost $60\%$ of these students pay for none of their assigned courses, with about $4\%$ obtaining multiple course seats with positive prices. These results are in line with the ones presented in Section 3 that show a larger percentage of Year $4$ students being indifferent between their allocation in the PMP mechanism and the RSD benchmark than students of any other year of study (see Table 5). In earlier years of study, a larger percentage of students pay for course seats. Under $45\%$ of first- through third-year students obtain all their course seats for free, with more than $10\%$ needing to pay for multiple courses in each year.
@@ D.2 Alternative Course Allocation Mechanisms
In turn, the Random Serial Dictatorship with course reserves (RSD) benchmark does not respect any priority structure. The cause of such violations is the presence of course reserves that can be interpreted as seat-specific priorities. Each of a course’s reserved seats has a priority order over students that places reserve-eligible students above others in the same year of study. Unreserved seats only prioritize students by seniority. Other mechanisms in the literature have seat-specific priorities. For example, Hafalir, Yenmez and Yildirim (2013) introduces to the school choice literature a version of the Deferred Acceptance mechanism, the Deferred Acceptance mechanism with minority reserves.
@@
The Deferred Acceptance (DA) with Minority Reserves with single and multiple tie-breakings. The Deferred Acceptance with minority reserves mechanism was proposed by Hafalir, Yenmez and Yildirim (2013) to address affirmative action policies in school choice. This mechanism is similar to the Deferred Acceptance mechanism, but it allows each school to have a number of reserved seats designated for affirmative action students. These seats give higher priority to minority students up to the point that minorities fill the reserves. We extend this mechanism to a many-to-many matching setting with indifferences. We consider two cases: where indifferences are broken with a single tie-breaking rule and where indifferences are broken with multiple tie-breaking rules, one for each course. [^36] We do not claim that the theoretical properties of Hafalir, Yenmez and Yildirim (2013) extend to many-to-many settings with indifferences. We leave the analysis of this important question for further research.
@@
Probabilistic Serial (PS) with Seniority and Reserves. The PS mechanism was proposed by Bogomolnaia and Moulin (2001) to randomly assign objects to agents in settings with single-unit demand for objects and single-unit supply of object types. Kojima (2009) extended the mechanism to settings with multi-unit demand for objects, and Budish et al. (2013) extended the mechanism to settings with multi-unit supply of objects. The undergraduate course allocation setting requires both multi-unit demand and multi-unit supply of courses. The only paper that considers both multi-unit demand and multi-unit supply of courses with preferences over individual courses is an unpublished paper by Pycia (2011). [^37] Nguyen, Peivandi and Vohra (2016) also consider multi-unit demand and multi-unit supply of objects but consider agent preferences over bundles of objects. They propose a version of a probabilistic serial mechanism called the Bundled Probabilistic Serial mechanism that has been successfully implemented at the Technical University of Munich by Bichler and Merting (2021). One of the versions of their probabilistic serial mechanism could also be applied to settings with agent preferences over individual objects. The setting of undergraduate course allocation is further complicated by the presence of student seniorities and course reserves. We define an extension of the probabilistic serial mechanism called Probabilistic Serial with Seniority and Reserves.
@@
Table D.2: Each student’s preferred mechanism by year of study for the alternative mechanisms.
Deferred Acceptance with minority reserves (single) Deferred Acceptance with minority reserves (multiple) Random Serial Dictatorship with actual course reserves
[…]
Year 1 47.3 (1.3) 2.2 (0.7) 50.2 (1.2) 8.4 (1.0) 46.4 (1.5) 28.7 (1.3)
Year 2 23.7 (1.3) 33.6 (1.6) 29.9 (1.3) 36.6 (1.5) 19.1 (1.1) 63.6 (1.3)
Year 3 6.5 (0.8) 42.7 (1.5) 14.0 (1.0) 45.4 (1.6) 4.4 (0.5) 75.5 (1.2)
Year 4 2.0 (0.4) 20.3 (1.5) 12.3 (0.8) 28.1 (1.5) 0.8 (0.3) 53.6 (1.3)
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
[…]
Year 1 88.4 (0.9) 3.4 (0.3) 75.8 (1.5) 4.4 (0.4) 54.2 (1.7) 14.7 (1.3)
Year 2 74.6 (1.2) 13.1 (0.8) 63.7 (1.4) 19.0 (1.1) 31.2 (2.0) 49.6 (1.5)
Year 3 34.6 (1.5) 42.9 (1.5) 30.4 (1.4) 36.5 (1.3) 18.2 (1.6) 57.8 (2.1)
Year 4 15.8 (0.9) 47.3 (1.6) 19.0 (1.0) 20.4 (1.0) 17.2 (1.0) 41.6 (1.4)
@@
Notes: This table reports the percentage of students who prefer a certain mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Results are averages across 100 runs with different random component draws.
@@
Simulation results. Our simulation results for the six alternative mechanisms are presented in Tables D.2D.4. Table D.2, parallel to Table 5 in the main text, reports the percentage of students who prefer the allocation of a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Table D.3, parallel to Table 7, presents information about the change in the standard deviation of student utility by year of study for the six alternative mechanisms compared to the Random Serial Dictatorship with optimal course reserves. Table D.4, parallel to Table 8, presents the shares of students who experience schedule envy towards students of the same priority or lower priority. Table D.5 presents the percentage of students who experience a priority violation. A student experiences a priority violation if she wants to enroll in a course being taken by a student of strictly lower priority. The PMP, DA-STB, and DA-MTB mechanisms studied in the main text have no priority violations.
@@
The DA with minority reserves (with single and multiple tie-breakings) result in a larger percentage of junior students (Years 1 and 2) who prefer these mechanisms ($29.9\%-47.3\%$), and a smaller percentage of senior students (Years 3 and 4) who prefer these mechanisms to the RSD benchmark ($15.6\%-25.1\%$) compared to the PMP, DA-STB, and DA-MTB mechanisms (see Table 5). These results are a direct consequence of the presence of course reserves, which ensures better access to course seats for junior students. Course reserves do not provide equal access to seats, leading to an increase in the standard deviation of student utility for junior students ($6.0\%-14.7\%$) compared to the RSD benchmark. We see similar underperformance in other measures of allocation fairness. In both mechanisms, greater than $20\%$ of students experience schedule envy toward a student of weakly lower priority. Furthermore, $40\%$ of students face a priority violation, indicating that many students can benefit from a course seat assigned to a student of lower priority.
@@
Table D.3: The change in the standard deviation of students’ utilities by year of study for the alternative mechanisms.
Deferred Acceptance with minority reserves (single) Deferred Acceptance with minority reserves (multiple) Random Serial Dictatorship with actual course reserves
Year 1 14.7% 13.3% 13.2%
Year 2 7.1% 6.0% 3.7%
Year 3 -11.6% -13.8% -18.2%
Year 4 -3.4% -5.5% -13.0%
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
Year 1 -10.4% -11.4% 9.4%
Year 2 -12.5% -14.5% -0.8%
Year 3 -20.8% -18.4% -22.0%
Year 4 -22.5% -8.5% -14.1%
@@
Notes: The change in the standard deviation of student utility by year of study for the six alternative mechanisms compared to the Random Serial Dictatorship with optimal course reserves. The results are based on $100$ runs with different random component draws.
@@
The RSD with actual course reserves mechanism performs quite poorly compared to the RSD benchmark. For second- through fourth-year students, more than $50\%$ of students prefer the outcome of the RSD mechanism with optimal course reserves over the actual course reserves. The mechanism underperforms in terms of allocation fairness as well, with $27.8\%$ of students preferring the schedule of a student with weakly lower priority in each course and $59\%$ of students with priority violations. Only first-year students benefit from using the university’s actual course reserves. These results justify why we consider the RSD mechanism with optimal course reserves as the benchmark mechanism rather than the RSD mechanism with actual course reserves. Using the actual course reserves presents an unfair comparison, as course reserves are typically set large at the beginning of the process and relaxed later, resulting in the allocation of courses to students that, to a large extent, is determined by an unstructured decentralized process that happens after the main allocation round in practice.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the alternative mechanisms. The first column provides information about students who experience no envy. The other columns show the percentage of students who experience schedule envy bounded by $1,...,5$ courses. Results are averages across $100$ runs with different random component draws.
The A-CEEI and A-CEEI with Kludgy Pricing mechanisms do not strictly enforce course priorities through prices, unlike the PMP mechanism. These relaxed restrictions result in more course seats available to junior students (Year 1 and 2), who overwhelmingly strictly prefer each mechanism to the RSD benchmark ($63.7\%-88.4\%$). The results reverse for senior students (Years 3 and 4), with more students favoring the benchmark. Both mechanisms reduce the standard deviation of student utility across all years of study ($8.5\%-22.5\%$). This decrease is substantially larger than those observed in Section 3 for the PMP mechanism, suggesting that the relaxed priority constraints allow students to receive more equal access to courses. Similarly, over $95\%$ of students experience no envy toward a student of weakly lower priority, with schedule envy bounded by a single course in each mechanism. However, without these constraints, priority violations are common in both mechanisms, with $61.1\%$ of students in the A-CEEI mechanism and $44.6\%$ of students in the A-CEEI with Kludgy Pricing mechanism whose schedule would benefit from a course seat taken by a student of strictly lower priority. Also, the average timing to calculate the allocation of the A-CEEI and A-CEEI with Kludgy Pricing mechanisms (2003 and 1894 seconds, respectively) is almost twice as much as that of the PMP mechanism (1112 seconds).
@@
The Probabilistic Serial with Seniority and Reserves mechanism is the only one of mechanisms we consider that produces a random assignment of course seats to students. This approach is more flexible, allowing the assignment of probability shares of course seats to students rather than full seats. The mechanism, however, delivers more beneficial outcomes only to students of Year $1$ (see Table D.2). More second- through fourth-year students strictly prefer their schedule in the RSD benchmark. This result is in contrast to the PMP mechanism, where the percentage of students who prefer the PMP mechanism is larger than the percentage of students who prefer the RSD benchmark for second- through fourth-year students. In terms of allocation fairness, the probabilistic serial mechanism reduces the standard deviation of students’ utilities for all years of study and results in just $10.9\%$ of students with envy towards a student of weakly lower priority. [^38] Kojima (2009) and Budish et al. (2013) establish that the probabilistic serial mechanism in multi-unit demand setting and multi-unit supply setting, respectively, results in no envy. This result does not extend to settings where agents demand multiple units of objects and objects are supplied in multiple units simultenously. However, in doing so, the mechanism results in a large percentage of students who experience priority violation ($84.6\%$). This can be explained, in part, by the flexibility of the mechanism. We count a student as experiencing a priority violation if there is a positive probability share of a course that she prefers to some assigned probability share of hers (possibly empty) that is instead assigned to some student at a strictly lower priority level. This matches the definition of ex-ante stability proposed in Kesten and Ünver (2015). We also consider envy in terms of ex-ante assignment, not an ex-post assignment, as for the other alternative mechanisms. Hence, the results are not directly comparable.

### v2

@@ C.4 Calibration
We use the Random Serial Dictatorship with adjusted course reserves to calculate simulated moments and find $\pi^{*}=(\theta^{*},\Gamma^{*},z^{*})$ that minimize a criterion function:
$m-m^{s}(\pi)||_{W}^{2}=(m-m^{s}(\pi))^{\prime}W(m-m^{s}(\pi)),$ (C5)
where $m$ is a set of moments constructed from the empirical data, $m^{s}(\pi)$ is the average of moments constructed from $100$ simulation rounds for a given vector $\pi=(\theta,\Gamma,z)$ (drawing random student utility components and choice sets for each round), and $W$ is a weight matrix. We use the following procedure (see Jalali, Rahmandad and Ghoddusi, 2015):
@@
  • Second Stage. We start with the optimal weight matrix $W^{*}_{\ell}=W^{*}$ for $\ell=1$. We then use the optimal weight matrix to obtain more precise estimates for parameters $\pi^{*}_{\ell}$. Using the more precise estimate $\pi^{*}_{\ell}$, we compute the optimal weight matrix $W^{*}_{\ell+1}$ and repeat the process $\ell=2,3,$... until the maximum change of calibrated parameters $|\pi^{*}_{l+1}-\pi^{*}_{l}|$ becomes smaller than $0.01$ to ensure convergence.
@@
The horizontal components are positive for students in their first and second years and mostly negative in their third and fourth years. This observation suggests that students prefer enrolling in more courses during their initial years of study. The tendency to enroll in fewer than five courses in later years may be due to several factors. Such students may choose part-term courses that are not in our data or have already accumulated enough credits for their degree, allowing them to ease their course load in the final years.
The middle portion of the table presents interaction terms of matrix $\Gamma$. The general trend of negative utility changes emphasizes a preference for internal college courses, but specific positive changes indicate potential areas for beneficial inter-college course offerings. For example, students from colleges $C$ and $D$ value highly courses in college $A$.
The bottom portion of the table presents quantiles of $756$ coefficients of vertical component $z$, with each coefficient $z_{c}$ being responsible for the average popularity of the course among students from the same college. The values range from a minimum of 2.32-2.32 to a maximum of $2.28$. This wide range indicates significant variation in the popularity of courses within colleges. Most of the quantile values are negative, up to the $90th$ percentile, indicating that a majority of the courses have negative values. This result is most likely manifested by the student utility model that relies on college-year parameters to identify student preferences. Students randomly draw $80$ courses in their choice set but take no more than $5$ courses. Our model, which relies on college-level and year-level parameters, matches this empirical pattern by putting the values of the vertical components of most courses in the negative region.
@@ Appendix D Additional Simulations
This appendix provides additional simulation results. First, for the simulations conducted in the main text, we present additional results on the equilibrium outcomes of the Pseudo-Market with Priorities (PMP) mechanism (Section D.1) and compare performance with some alternative mechanisms (Section D.2). Second, we supplement the main text with simulation results with a department-first priority structure (Section D.3). Last, we report robustness checks by considering several sizes for student choice sets and the standard deviation of the noise utility parameter (Section D.4).
D.1 Outcomes of the Pseudo-Market with Priorities Mechanism
First, we present statistics on the equilibrium prices in the PMP mechanism for the simulations reported in Section 3.3. The top part of Table D.1 provides information about where the cutoff levels of priority are across courses and how high the prices are for students at the cutoff level of priority. The bottom part provides information about the percentage of students within each year of study allocated different numbers of courses with positive prices.
Around $70.8\%$ of courses have their cutoff priority at level $1$. The large share of courses at the lowest priority level reflects the non-binding enrollment of most courses in the university data, where just $11.2\%$ of courses are at or above their maximum enrollment capacity (see Table 2 on p. 2). Above level $1$, there is a noticeable difference in the percentage of courses with odd and even cutoff levels of priority. Substantially more courses have odd cutoff levels ($8.0\%$, $8.5\%$, $5.9\%$ for levels $3$, 55, and $7$, respectively) than even ones ($2.1\%$, $2.6\%$, $1.7\%$, $0.5\%$ for levels $2$, 44, 66, and $8$). The hybrid priority structure we utilize places students at odd priority levels in a course if they are ineligible for any of its reserved seats. The students eligible for a course’s reserved seats typically only comprise a small subset of the interested students. Furthermore, some courses do not reserve any seats. As a result, the cutoff level of priority often ends up at an odd level rather than an even one.
@@ D.1 Outcomes of the Pseudo-Market with Priorities Mechanism
Cutoff level of priority 11 22 33 44 55 66 77 88
@@
Notes: This table presents statistics on the equilibrium prices in the Pseudo-Market with Priorities mechanism. Results are averages across $100$ runs with different random component draws.
Among courses with a cutoff priority level of $1$, the average price at the cutoff is just $0.06$. As many courses are free for all students, the average cutoff price for these courses is considerably lower than for courses with a higher cutoff level. For higher cutoff levels, the average cutoff prices lie between $0.75$ and $0.94$, with higher prices for odd cutoff levels. Student budgets are assigned in $[1,1.25]$, meaning that, at the average cutoff price, each student can afford one but not two courses for which they are at the cutoff level of priority. Still, students manage to take $k=5$ courses by taking courses for which they are above the cutoff level of priority, where they face a priority-specific price of 0.
The bottom part of Table D.1 further explores the number of courses students pay a positive price for. Across all years of study, at least $36\%$ of students obtain all their course seats for free, and fewer than $13\%$ of students pay a positive price for more than one course. Many courses have a zero price for all students, leading some students to obtain their most preferred course schedule for free. In addition, students often cannot afford a second course with a positive price after obtaining one.
These findings differ across years of study. Less than $7\%$ of cutoff levels of priority are $7$ or $8$, implying that Year $4$ students face zero price for the vast majority of courses. Almost $60\%$ of these students pay for none of their assigned courses, with about $4\%$ obtaining multiple course seats with positive prices.
In earlier years of study, a larger percentage of students pay for course seats. Under $45\%$ of first- through third-year students obtain all their course seats for free, with more than $10\%$ needing to pay for multiple courses in each year.
Table D.2: Over-enrolled courses in the Pseudo-Market with Priorities mechanism.
# seats above max capacity $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses $2.5$ $0.4$ $0.1$ $0.02$ $0.0$
We also present results on course over-enrollment in the PMP mechanism’s outcomes. The average market-clearing error in our simulations is $21.4$, less than half of the theoretical worst-case bound of $\alpha\approx 43.5$. Table D.2 reports that $2.5\%$, or roughly 19 courses, are over-enrolled by one seat. In addition, $0.4\%$ (3 courses) are over-enrolled by two seats, and $0.12\%$ (<<1 course) are over-enrolled by at least three seats. No courses are over-enrolled by more than four seats. In contrast, the university’s original allocation has $7.3\%$ of courses over-enrolled, with $1.9\%$ over-enrolled by at least five seats (see Table 2). A lower market-clearing error and over-enrollment are possible by allowing a longer runtime for our algorithm.
@@ D.2 Alternative Course Allocation Mechanisms
In turn, the Random Serial Dictatorship with course reserves (RSD) benchmark does not respect any priority structure. The cause of such violations is the presence of course reserves that can be interpreted as seat-specific priorities. Each of a course’s reserved seats has a priority order over students that places reserve-eligible students above others in the same year of study. Unreserved seats only prioritize students by year of study. Other mechanisms in the literature have seat-specific priorities. For example, the Deferred Acceptance with minority reserves mechanism introduced in Hafalir, Yenmez and Yildirim (2013) is a modification of the Deferred Acceptance mechanism with seat-specific priorities.
@@
Deferred Acceptance (DA) with Minority Reserves with single and multiple tie-breakings. The Deferred Acceptance with minority reserves mechanism was proposed by Hafalir, Yenmez and Yildirim (2013) to address affirmative action policies in school choice. This mechanism is similar to the Deferred Acceptance mechanism, but it allows each school to have a number of reserved seats designated for affirmative action students. These seats give higher priority to minority students up to the point that minorities fill the reserves. We extend this mechanism to a many-to-many matching setting with indifferences, using the optimal course reserves discussed in Section 3.2. We consider two cases: where indifferences are broken with a single tie-breaking rule and where indifferences are broken with multiple tie-breaking rules, one for each course. [^33] We do not claim that the theoretical properties of Hafalir, Yenmez and Yildirim (2013) extend to many-to-many settings with indifferences. We leave the analysis of this important question for further research.
@@
Probabilistic Serial (PS) with Seniority and Reserves. The PS mechanism was proposed by Bogomolnaia and Moulin (2001) to randomly assign objects to agents in settings with single-unit demand for objects and single-unit supply of object types. Kojima (2009) extended the mechanism to settings with multi-unit demand for objects, and Budish et al. (2013) extended the mechanism to settings with multi-unit supply of objects. The undergraduate course allocation setting requires both multi-unit demand and multi-unit supply of courses. The only paper that considers both multi-unit demand and multi-unit supply of courses with preferences over individual courses is an unpublished paper by Pycia (2011). [^34] Nguyen, Peivandi and Vohra (2016) also consider multi-unit demand and multi-unit supply of objects but consider agent preferences over bundles of objects. They propose a version of a probabilistic serial mechanism called the Bundled Probabilistic Serial mechanism that has been successfully implemented at the Technical University of Munich by Bichler and Merting (2021). One of the versions of their probabilistic serial mechanism could also be applied to settings with agent preferences over individual objects. The setting of undergraduate course allocation is further complicated by the presence of student seniorities and course reserves. We define an extension of the probabilistic serial mechanism called Probabilistic Serial with Seniority and Reserves, using the optimal course reserves discussed in Section 3.2.
@@
Figure D.1: Each student’s preferred mechanism for the alternative mechanisms.
@@
Notes: This figure reports the percent of students who strictly prefer each alternative mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark. Results are averages across 100 runs with different random component draws.
@@
Simulation results. Our simulation results for the six alternative mechanisms are presented below. Figure D.1, as in Figure 1 in the main text, reports the percentage of students who prefer the allocation of a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Figure D.2, as in Figure 2, reports the percent improvement in mean utility among students with changing schedules for the six mechanisms. [^35] On average, the DA-STB with minority reserves mechanism changes schedules for 775, 923, 698, and 319 students in Years 1, 2, 3, and 4, respectively; the DA-MTB with minority reserves mechanism does so for 918, 1071, 845, 576 students; the RSD mechanism with actual course reserves does so for 1176, 1131, 1135, and 774 students; the A-CEEI mechanism does so for 1437, 1413, 1102, and 898 students; the A-CEEI with kludgy pricing mechanism does so for 1256, 1131, 951, and 562 students; and the PS mechanism does so for 1078, 1302, 1080, and 838 students. Figure D.3, as in Figure 3, presents the change in the standard deviation of student utility for each mechanism by year of study. Table D.3, parallel to Table 5, reports students who experience schedule envy towards students of the same priority or lower priority. Table D.4 presents the percentage of students who experience a priority violation: who want to enroll in a course being taken by a student of strictly lower priority. The PMP, DA-STB, and DA-MTB mechanisms studied in the main text each have no priority violations.
Figure D.2: Improvements in utility among students with changing schedules for the alternative mechanisms.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each alternative mechanism over the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
@@
The DA with minority reserves (with single and multiple tie-breakings) mechanisms each have a smaller percentage of students who prefer the mechanism than students who prefer the benchmark. This negative result contrasts with the result for their deferred acceptance counterparts, as studied in Section 3. These mechanisms only benefit Year 1 students, who significantly improve in mean utility ($39.1\%-46.9\%$ among students with changing schedules). These results are a direct consequence of the presence of course reserves, which ensure better access to course seats for junior students. Course reserves do not provide equal access to seats, leading to an increase in the standard deviation of student utility for junior students ($6.0\%-14.7\%$) compared to the RSD benchmark. We see similar underperformance in other measures of allocation fairness. In both mechanisms, more than $20\%$ of students experience schedule envy toward a student of weakly lower priority. Furthermore, more than $40\%$ of students face a priority violation, indicating that many students can benefit from a course seat assigned to a student of lower priority.
@@
Figure D.3: The change in the standard deviation of students’ utilities for the alternative mechanisms.
@@
Notes: This figure reports the percent change in the standard deviation of students’ utilities for each alternative mechanism compared to the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
@@
The RSD with actual course reserves mechanism performs quite poorly compared to the benchmark. More than $50\%$ of students strictly prefer the RSD mechanism with optimal course reserves over the RSD mechanism with actual course reserves. Using the actual course reserves only improves outcomes for first-year students, who have greater access to reserved seats. The mechanism underperforms in terms of allocation fairness as well, with $27.8\%$ of students preferring the schedule of a student with a weakly lower priority in each course and with $59\%$ of students experiencing priority violations. These results justify why we consider the RSD mechanism with optimal course reserves as the benchmark mechanism rather than the RSD mechanism with actual course reserves. Using the actual course reserves presents an unfair comparison, as the university sets reserves large at the beginning of the process and relaxes them later. This results in an allocation of courses to students, which, to a large extent, happens after the main registration process in practice.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the alternative mechanisms. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across 100 runs with different random component draws.
The A-CEEI and A-CEEI with Kludgy Pricing mechanisms do not strictly enforce course priorities through prices, unlike the PMP mechanism. These relaxed restrictions result in more course seats available to junior students (Years 1 and 2), who overwhelmingly prefer each mechanism to the RSD benchmark (see the bottom section of each bar in Figure D.1). While the results are reversed for senior students (Years 3 and 4), many more students overall prefer each of these mechanisms ($48.4\%-54.8\%$) than the number of students who prefer RSD ($19.7\%-25.7\%$). Both mechanisms reduce the standard deviation of student utility across all years of study ($8.5\%-22.5\%$). This decrease is substantially larger than that observed in Section 3 for the PMP mechanism, suggesting that the relaxed priority constraints allow students to receive more equal access to courses. Similarly, over $95\%$ of students experience no envy toward a student of weakly lower priority, with schedule envy bounded by a single course in each mechanism. However, without these constraints, priority violations are common in both mechanisms, with $61.1\%$ of students in the A-CEEI mechanism and $44.6\%$ of students in the A-CEEI with Kludgy Pricing mechanism whose schedule would benefit from a course seat taken by a student of strictly lower priority. Additionally, the average time to calculate the allocation for the A-CEEI and A-CEEI with Kludgy Pricing mechanisms (2003 and 1894 seconds, respectively) is almost twice that of the PMP mechanism (1112 seconds).
@@
The PS mechanism is the only mechanism considered here that produces random assignments. Random mechanisms are more flexible, allowing the assignment of probability shares of course seats to students rather than full seats. The PS mechanism, however, delivers more beneficial outcomes only to Year $1$ students. More second- through fourth-year students prefer their schedule in the RSD benchmark than the number of students who prefer their PS outcome. These outcomes are also reflected in utility improvements among students with changing schedules. There is a $31.8\%$ improvement in average utility for Year 1 students. In contrast, all the other student cohorts experience a decrease in average utility (see Figure D.2). In terms of allocation fairness, the probabilistic serial mechanism reduces the standard deviation of students’ utilities for all years of study and results in just $10.9\%$ of students with envy towards a student of weakly lower priority. [^36] Kojima (2009) and Budish et al. (2013) prove that the probabilistic serial mechanism results in no envy in multi-unit demand and multi-unit supply settings, respectively. This result does not extend to settings where agents demand multiple units of objects and objects are supplied in multiple units simultaneously. In doing so, the mechanism results in a large percentage of students who experience priority violation ($84.6\%$). This can be explained, in part, by the flexibility of the mechanism. We count a student as experiencing a priority violation if there is a positive probability share of a course that she prefers to some assigned (possibly empty) probability share that is instead assigned to some student at a strictly lower priority level. This matches the definition of ex-ante stability proposed in Kesten and Ünver (2015). We also consider envy in terms of the ex-ante assignment, not an ex-post assignment. Hence, the results are not directly comparable.